A robot robe drops a camera off the rim of 239 m high cliff on Mars, where the free-fall acceleration is -3.7m/s^2
A. Find the velocity with which the camera hits the ground.
B. Find the time required for It to hit the ground.
Answer #1

a)

using 3rd equation of motion

v^2 = 2 g h

v^2 = 2 * 3.7* 239

v = 42.05 m/s

======

b)

t = v/g = 42.05 / 3.7 = 11.366 s

======

Comment before rate in case any doubt, will reply for sure.. goodluck